### v1

# Implementing LLMs in Industrial Process modeling: Addressing categorical variables

Abstract
Important variables of processes are, in many occasions, categorical, i.e. names or labels representing, e.g. categories of inputs, or types of reactors or a sequence of steps. In this work, we use Large Language Models (LLMs) to derive embeddings of such inputs that represent their actual meaning, or reflect the “distances" between categories, i.e. how similar or dissimilar they are. This is a marked difference from the current standard practice of using binary, or one-hot encoding to replace categorical variables with sequences of ones and zeros. Combined with dimensionality reduction techniques, either linear such as Principal Components Analysis (PCA), or nonlinear such as Uniform Manifold Approximation and Projection (UMAP), the proposed approach leads to a meaningful, low-dimensional feature space. The significance of obtaining meaningful embeddings is illustrated in the context of an industrial coating process for cutting tools that includes both numerical and categorical inputs. The proposed approach enables feature importance which is a marked improvement compared to the current state-of-the-art (SotA) in the encoding of categorical variables.
1 Introduction
Machine learning is currently extensively used in process modeling, design and engineering for a wide array of tasks, ranging from prediction with regression methods [1, 2], clustering and classification [3, 4, 5], the development of surrogate models of digital twins [6, 7, 8] and optimization and control [9, 10].
When dealing with industrial/production data, it is often the case that data are not numerical values, e.g. temperature time-series, flow-rates, pressure etc., but rather involve categorical variables, i.e. categories described by an assigned name. Examples include serial numbers of reactors, working-names of products or even an entire sequence of steps represented by a single name. One of the most popular methods to address this issue is one-hot encoding, which replaces categorical variables with several columns of dummy variables, commonly a sequence of zeros and ones. By replacing categorical variables with numerical ones, it is possible to use them as features in various machine learning algorithms. However, by doing so, the resulting representation is agnostic of the actual meaning of the variables, and it is unable to reflect the actual "distance" (i.e. relative similarity or dissimilarity) between the different categories. For these reasons, it is an obstacle for feature importance and sensitivity analysis, uncertainty quantification, and/or optimization, diminishing the opportunity to create insights from the data and increase understanding of the physical process at hand.
Following their rise to popularity, a lot of interest has fallen in LLMs and their implementation [11, 12]. Specifically in chemical and process engineering, the prospects of LLMs for education [13] and process mining [14] have been examined. With their work, Decardi-Nelson et al. [15] investigate the implementation of the generative ability of LLMs in the various stages of process systems engineering (PSE), namely synthesis and design, optimization and integration, and process monitoring and control. In [16], the use of an agent-based LLM approach in the context of process simulation is proposed, combining four different LLM-based agents for the observation of the process. Furthermore [17] evaluate the capabilities of LLMs for the prediction, generation and understanding of tabular data. However, despite the widespread use of LLMs for text embedding, exemplified by S-Bert [18], to the best of our knowledge, the impact of using LLMs to encode categorical variables in the development of predictive models has not been explored.
With this study, we aim to explore the potential benefits of implementing LLMs to derive meaningful embeddings of categorical data, using as inputs short textual descriptions of what the different categories represent. These embeddings are then compressed using either linear or nonlinear dimensionality reduction methods and used as inputs to tree-based regression models. Actual production data are used in this study, stemming from an industrial chemical vapor deposition reactor used for the production of wear-resisting coatings on cutting tools, called “inserts". This process combines competing physical and chemical mechanisms in very complex and ever-changing geometrical set-ups. There are no inline measurements and therefore there is no possibility to monitor the process as it happens and intervene with control actions. Coating thickness measurements are collected ex-situ, after the process is finished, providing information about the quality of the product.
Aiming to develop a data-driven predictive model [19] for the quality characteristics of the coating, a tree-based regression algorithm is trained using several numerical and categorical inputs. The categorical inputs include the insert (cutting tool) name, a series of letters and numbers, representing an object with an ISO-specified shape and size. In what we call the “original" (SotA) treatment of the data, the entire name is encoded with one-hot encoding, which is oblivious of the meaning of each component of the name. This poses a restriction on the amount of information that may be extracted from the data-driven model. As an example, it is not possible to establish if and what effect the distribution of inserts in the reactors, their placement, in general, has on the output, i.e. the quality of the product.
In this work, several models, pre-trained an un-trained, are compared and subsequent feature importance analysis, using Shapley values, reveals that critical process inputs; physically meaningful and useful conclusions arise when using the proposed embedings in contrast to binary encoding.
The development of the proposed tool leads to improved understanding of the process set up since it paves the way for sensitivity analysis and uncertainty quantification. Furthermore, it allows the incorporation into datasets of pieces of production information that is maintained in the form of notes, short texts, in the production diaries. This information is often very important and is assessed by experts in the production line but not in a systematic way.
2 Process description
The present work uses the example of a two-step coating process, which takes place in a commercial Chemical Vapor Deposition (CVD) reactor (Sucotec SCT600TH). The process is described in detail in previous works, which the interested reader can refer to [19, 20, 21] and is summarized here for completeness. First, a 9 m Ti(C,N) layer is deposited on the cemented carbide cutting inserts, examples of which are presented in Fig. 1(a). This step is followed by the deposition of an alumina layer under a AlCl3–CO2–HCl–H2–H2S chemical system, with the process temperature and pressure for the -Al2O3 deposition being =1005°C and =80 mbar, respectively [22].
The CVD reactor consists of 40-50 perforated disks stacked vertically, each accommodating inserts. Fig. 1(b) provides a schematic 3D illustration of three such disks for clarity. Gas reactants enter the reactor through perforations in a rotating cylindrical tube positioned at the center of the disk stack. Each disk level features two diametrically opposite perforations, with a 60° angular offset between the axes connecting the inlet holes at each level. The rotating motion of the centrally located inlet tube, at a speed of 2 RPM, makes the process inherently periodic. An important aspect of the process is the fact that the internal geometry of the reactor varies between production runs as the geometry of the inserts and the disks on which they are placed is modified according to production requirements.
The primary objective of the process is to achieve a uniform coating thickness, as this uniformity directly correlates with consistent product longevity [23]. Ideally, this uniformity would be consistent across all production runs, reactors, and sites. However, this is not always achieved. Therefore, it is crucial to develop a method for predicting the coating thickness of the inserts based on the reactor setup. In addition, establishing a systematic approach to evaluate factors affecting the uniformity of the coating thickness is essential. To this extent, the application of both equation-based methods [20] and data-driven methods [21, 19] has been demonstrated in previous work, to which the interested reader is referred for further information on the process.
3 The original data-driven approach
At each production run, 15 thickness measurements are taken ex-situ, using the Calotest method [24]. A 2D representation of the reactor indicating the points where thickness is typically measured is shown in Fig. 2. It should be noted though that, due to production decisions, additional measurements are taken in the R position (the one closest to the reactor outlet). For this reason, the focus of the -Al2O3 coating thickness predictions falls on the R position of the reactor.
In addition to coating thickness measurements, the dataset also contains several features regarding the set-up of the production run. These features are used as inputs for our predictive ML models. An important feature is the production “recipe”, which encapsulates the steps taken and the process conditions during production. Furthermore, there are features that contain information regarding the “how” and the “where” the inserts are placed within the reactor. More specifically, in the original approach, the following features are used: a) The number of inserts placed on each disk. b) The position of each disk within the reactor. c) The type of insert placed on each disk. Each type of insert has different geometrical characteristics. d) The surface area of the inserts placed on each disk.
It must be noted here that the categorical feature “Insert geometry" corresponds to ISO designations for indexable inserts [25], which are codes composed of eight alphanumeric characters. Those characters represent some measurements (numerical features) or even textual descriptions (categorical features) related to the insert geometry, as shown in Fig. 3.
Additional features are engineered which include the total surface area and the standard deviation of the surface area of the to-be coated inserts. The information available for the neighboring disks, i.e. the disks above and below the disk of interest, are also used for the development of our predictive models. Subject matter expertise suggests that the difference between the nominal surface area indicated in the production “recipe” and the actual surface area of the inserts within the reactor, is an important feature.
Following this step, ten features, both numerical and categorical, are available for the development of the predictive model, as presented in Table 1. The numerical features are standardized: centered (subtraction of the mean) and scaled (divided by the standard deviation), while the categorical variables are encoded using binary encoding [26].
| Feature | Type | Pre-processing | ||
| Number of inserts on disk | Numerical (integer) | standardization | ||
| Surface area of inserts on disk | Numerical (float) | standardization | ||
| Disk position | Numerical (integer) | standardization | ||
Numerical (float) | standardization | ||
| Surface area standard deviation | Numerical (float) | standardization | ||
Numerical (float) | standardization | ||
| Production “recipe” | Categorical | binary encoding | ||
| Insert geometry | Categorical | binary encoding | ||
| Insert geometry – disk above | Categorical | binary encoding | ||
| Insert geometry – disk below | Categorical | binary encoding |
4 Methods
This section provides a succinct description of the methods and algorithms used to create vector representations of insert shape descriptions through various embeddings, and to present similarities between the different representations. These vectors are included as numerical features in a regression model that predicts insert thickness based on other insert features listed in Section 3. Additionally, the model performance will be evaluated and compared across different embeddings using a cross-validation method, followed by feature importance and Shapley analysis.
4.1 Embeddings and transformers
Embedding is a method used to convert words or sentences from a vocabulary into dense numerical vectors suitable for machine learning models [27, 28]. More precisely, each word or sentence is mapped to a vector in Euclidean space, known as the embedded latent space. This representation helps capture and understand the semantic similarity between words.
In addition, transformers are a type of neural network architecture [29] that include an embedding layer, which encodes semantically similar words or sentences as close elements in the latent space [30]. Then, these deep learning methods rely on attention mechanisms to process sequential data. In particular, transformers are designed to manage context more effectively when processing text, allowing each word in a sentence to connect with every other word. This ability is essential for understanding meaning within context.
That said, we will now focus on the models used here: Doc2Vec and two pre-trained models (all-MiniLM-L12-v2 and all-mpnet-base-v2), respectively, whose main characteristics are summarized in Table 2.
Model Dimension Pre-trained Model Training Data Doc2Vec set to 3 no available insert shape descriptions all-MiniLM-L12-v2 384 yes 1B+ training pairs all-mpnet-base-v2 768 yes 1B+ training pairs
Both transformers were taken from Hugging Face Hub, which contains a vast collection of public Sentence Transformers models, extensively evaluated for their ability to embed sentences. The all-mpnet-base-v2 model provides the best quality and embedding performance, while all-MiniLM-L12-v2 is three times faster and still offers good embedding quality. It is also the best-performing model among those with a smaller latent space.
4.1.1 Doc2Vec
Doc2Vec [31] [32] is an unsupervised machine learning algorithm from the Gensim library to obtain vector representations of documents. In this case, a document is an object of the text sequence type and it could be anything from a short character string to documents structured by paragraphs such as an article or a book.
This algorithm is an extension of the Word2Vec, based on the bag-of-words [28] and bag-of-n-grams [28, 33]. It was designed to create fixed-length vector representations from pieces of texts. This model allows us to capture semantic meanings (distances between the words) and context (relationships between words and documents), which increase the accuracy of various tasks such as document classification, clustering, and recommendation.
Initially, we pre-process the documents to ensure agreement with the Doc2Vec model. This preprocessing includes the removal of punctuation and special characters, tokenization (word by word), and the creation of a dictionary. These steps are crucial to standardize the text and reduce noise, thereby improving the quality of the data and the resulting vectors.
The training process involves some parameters with which we can experiment to evaluate the model performance.
vector_size: the dimensionality of the feature vectors.
min_count: words that appear less than times are excluded from the training to focus on more frequent terms.
epochs: the model is trained for epochs to ensure adequate learning of patterns in the text.
4.1.2 all-MiniLM-L12-v2
The all-MiniLM-L12-v2 [6, 34] is a compact and efficient pre-trained language model designed for a variety of natural language processing (NLP) tasks [35]. It has been fine-tuned with more than 1 billion textual pairs to balance computational efficiency with high performance [36].
In addition, it utilizes a 12-layer deep network with attention mechanisms that allow to capture contextual relationships within text sequences efficiently. This understanding is essential for text classification, sentiment analysis, and named entity recognition tasks.
Pre-processing techniques also involve tokenization, which breaks down input text into meaningful units for the model to process, and data augmentation strategies to increase the model’s ability to generalize [37]. Then, the model produces embeddings of fixed length equal to 384, which are included in the training of our predictive model.
4.1.3 all-mpnet-base-v2
The all-mpnet-base-v2 model [38, 34] is a pre-trained transformer-based architecture developed by Microsoft. The all-mpnet-base-v2 model operates with a transformer encoder comprising 12 layers, each with 12 attention heads, and a hidden size of 768. It has been fine-tuned with more than 1 billion textual pairs to capture a wide array of linguistic patterns and contextual representations, making it highly suitable for various NLP tasks [35, 36].
Pre-processing involves tokenizing the input text, which is then fed into the model to produce embeddings of length 768. These embeddings are utilized in our experiments to explore various aspects of language understanding and contextual relevance.
4.1.4 Similarity between descriptions
Cosine similarity is a widely used metric in NLP for assessing the similarity between two vectors. In the context of word embeddings, cosine similarity is a valuable method for comparing the similarity of words or documents [39, 40].
It measures the cosine of the angle between two vectors in a high-dimensional space. The similarity score ranges from -1 to 1, where a score of -1 signifies completely opposite vectors, 0 indicates no similarity, and 1 means identical vectors. The cosine similarity between two vectors is computed using the following formula:
where represents the dot product between vectors and , while denotes the Euclidean norm.
4.2 XGBOOST
Following previous work [21], where an XGBOOST (Extreme Gradient Boosting) regressor [41] demonstrated excellent performance, XGBOOST is also used in the present implementation. XGBOOST is an ensemble [42] tree-based method [43] that allows both classification and regression. The XGBOOST algorithm builds shallow trees sequentially, with each tree being fit using the error residuals (difference between actual and predicted values) of the previous model. For more information on XGBOOST, the reader is referred to the original paper by Chen and Guestrin [41].
4.2.1 K-fold cross-validation
To ensure reliable predictions and avoid overfitting, it is common practice to reserve a portion of the data as a test set. However, if the test set is used for other purposes, such as fine-tuning hyperparameters, a validation set is also needed: the model is trained on the training set, validated on the validation set, and finally evaluated on the test set [40].
However, splitting the data into three sets can reduce the amount of data available for training. The -fold cross-validation [44] manages this by dividing the data into -folds. The model is trained on folds and tested on the remaining fold. This process is repeated times, and the final performance measure is the average across all folds.
Therefore, this approach is more efficient than using a fixed validation set, as it reduces computational cost and makes better use of the available data.
4.2.2 Feature Importance
Feature importance is a method used to measure the significance of input features in predicting a target variable by assigning them a score based on their impact [45]. Various techniques can be employed to determine feature importance, including statistical correlation measures [46], coefficients from linear models [47], and decision tree-based methods [47].
In this study, we utilize the XGBOOST model to analyze feature importance [48]. Thus, after constructing the boosted trees, it provides a way to retrieve the importance scores for each feature. In particular, we will use the total_gain score, which represents the total contribution of a feature to the model, calculated by considering the contribution of each feature across all trees [49, 50]. A higher total_gain value compared to other features indicates that the feature plays a more significant role in generating predictions.
4.2.3 SHAP Analysis
Shapley values, introduced by Shapley [51] and later applied to machine learning models in [52, 53], assess the average contribution of each feature to predictions. This helps us to understand how changes to a variable affect the model’s output. The core concept behind Shapley value-based explanations in machine learning is to allocate credit for a model’s output fairly among its input features, based on cooperative game theory principles. In this framework, the input features are treated as players in a game, where each player can either participate or not. If a player (input feature) joins, its value is known; if not, its value remains unknown. Shapley values are additive, meaning that in the context of model explanation, the sum of all SHAP values across the input features will always equal the difference between the baseline (expected) output and the actual output for the prediction being explained.
In this work, the SHAP (SHapley Additive exPlanations) analysis is applied to the proposed XGBOOST regression models. The resulting Shapley values will highlight the importance and influence of each feature on the model output.
Finally, to obtain an aggregated average score of the Shapley values, which allows us to derive the average contributions of each feature, we first sum the rows corresponding to each feature. Next, we take the absolute value of the resulting vector to ensure a positive vector of individual contributions. Then, we compute the average of these contributions.
5 Results
5.1 Data preprocessing
Regarding the insert names, each ISO designation code (cf. Fig. 3) is broken down into its various features, aiming to provide additional information that could be useful to include in the XGBOOST model.
Feature Type Pre-processing Number of inserts on disk Numerical (integer) standardization Surface area of inserts on disk Numerical (float) standardization Disk position Numerical (integer) standardization Total surface area of inserts inside the reactor Numerical (float) standardization Surface area standard deviation Numerical (float) standardization Nominal “recipe” surface area – actual surface area Numerical (float) standardization Insert clearance angle Numerical (float) standardization Insert clearance angle – disk above Numerical (float) standardization Insert clearance angle – disk below Numerical (float) standardization Insert cutting length Numerical (float) standardization Insert cutting length – disk above Numerical (float) standardization Insert cutting length – disk below Numerical (float) standardization Insert thickness Numerical (float) standardization Insert thickness – disk above Numerical (float) standardization Insert thickness – disk below Numerical (float) standardization Insert cutting length tolerance Numerical (float) standardization Insert cutting length tolerance – disk above Numerical (float) standardization Insert cutting length tolerance – disk below Numerical (float) standardization Insert thickness tolerance Numerical (float) standardization Insert thickness tolerance – disk above Numerical (float) standardization Insert thickness tolerance – disk below Numerical (float) standardization Production “recipe” Categorical binary encoding Insert shape Categorical Embedding* + standardization Insert shape – disk above Categorical Embedding* + standardization Insert shape – disk below Categorical Embedding* + standardization Insert characteristic Categorical binary encoding Insert characteristic – disk above Categorical binary encoding Insert characteristic – disk below Categorical binary encoding
The new features are summarized in Table 3). In the data transformation stage, we first apply data imputation to ensure that missing values were not included in our analysis. Then, we standardize the numerical features and use embeddings/transformers to represent categorical features as vectors.
Subsequently, the Doc2Vec model is trained for 2000 epochs. The vocabulary used for training consists of at least twelve short descriptions of insert shapes distributed across the current disk position and the positions above and below it. Therefore, we set the dimension of the dense vectors to the floor of the average number of words used in each description: vector_size = 3, and set the min_count parameter to 1 due to the short length of these descriptions.
In contrast, for the two pre-trained transformers, we use their most recent versions, which were trained on over one billion text pairs. Moreover, the resulting dense vectors have fixed and predetermined sizes: 384 for the all-MiniLM-L12-v2 model and 768 for the all-mpnet-base-v2 model.
5.2 Similarity
The first step is to verify whether the vectors that represent the embeddings by the different models are truly capturing the meaning of the sentences, including both semantic and contextual. Recall that each sentence has been mapped as vectors in Euclidean space. Consequently, we can consider measures such as the distance between these vectors or the angle between two vectors, as indicators of sentence similarity (see Section 4.1.4).
In this section, we propose to study the values obtained by calculating the cosine_similarity between each pair of dense vectors associated with the insert shape descriptions implemented in this study. These values are then collected and presented as heatmaps for each embedding and transformer.
Concretely, for the Doc2Vec model (see Fig. 4), we first observe a sort of reflexive property between descriptions, meaning each description is perfectly similar to itself. Additionally, we can see certain groups of descriptions that are expected to be similar, such as the group of rhombus-shaped inserts with different degrees. We also notice specific connections between certain shapes, like triangular and circular shapes with the rounded triangular shape, or some similarity among standard shapes like triangles, circles, and squares. It is also interesting to note that the filling material is close to the unknown inserts shape. This illustrates that the Doc2Vec model is capturing reasonably well the semantic and contextual meaning of the descriptions.
Additionally, in the case of the all-MiniLM-L12-v2 model, we again observe the reflexivity property. We also see similar patterns to those of the previous case, but more marked (refer to Fig. 5), with values surpassing 0.83, among semantically close insert shape descriptions, as expected. For instance, cosine_similarity values for rhombus or triangular shapes. Two additional observations stand out: the corresponding similarity values for the descriptions filling material and unknown shape with respect to other descriptions are below 0.44 in the first case and closer to zero in the second case. This indicates that the model is correctly capturing their meanings, and we have a reasonably accurate representation of the considered descriptions.
For the all-mpnet-base-v2 model, we obtain results similar to those of the all-MiniLM-L12-v2 model, maintaining the same patterns (see Fig. 6). Once again, the values for descriptions such as filling material or unknown shape are distinct from other groups, with values below 0.41. Additionally, there is a high similarity among the descriptions in the group of quadrilaterals located in the upper left of Fig. 6, with values exceeding 0.88. However, there is no clear differentiation between standard shapes such as squares, circles, or triangles, which have values in the middle and lower right part of this Fig. 6.
Now, it is worth noting that one of the advantages of the Doc2Vec model is that, unlike pre-trained transformers, we can vary the dimensionality of the latent space according to our needs. We first compare the implementation of Doc2Vec using latent spaces of larger size, specifically we select 384 and 768, i.e. the dimension of the embedding of the two pre-trained models, to the implementation of Doc2Vec with a three-dimensional embedding.
Thus, two more possible scenarios are interesting for further study of cosine_similarity at this point: the first regarding Doc2Vec, for which we can consider latent spaces with dimensions of 384 and 768, as in the case of both transformers. The results obtained are similar to those for a fixed dimensionality of 3.
In contrast, the second scenario corresponds to the high dimensionality of the latent spaces produced by both transformers, possible consequences of reducing their dimensionality can be explored. For this, we consider a linear method PCA of dimension 3. The results obtained are presented in Appendix A and show the same patterns observed when dimensionality reduction is not applied and capturing the same semantic and contextual meaning, as shown in Figs. 13 and 14. In the next section, we will also analyze how both cases impact the model’s performance.
5.3 XGBOOST model performance
In all cases, the vectors obtained by the transformers are included as numerical features in the XGBOOST model. At this point, it is interesting to explore how the amount of information considered during training influences the model’s performance. To address this, we will examine two important metrics in supervised machine learning models that provide insight into their accuracy and robustness: mean squared error (MSE) and score.
With this in mind, the next two sections will discuss the results obtained from applying the -fold cross-validation method with , while considering progressively increasing portions of the data used for training.
5.4 Mean squared error, MSE
In this section, we focus on the first metric, MSE. Results are presented as 99.9 confidence intervals for the average MSE, using a bandwidth of one standard deviation (Fig. 7). This follows from the fact that the MSE values have a normal distribution (null hypothesis), as shown using the Kolmogorov-Smirnov test in Table 4; the p-values (> 0.05), which represent the probability that the observed values occur under the null hypothesis, in each case allow us to assume the normality of the MSE values.
| Fraction |
| of data (%) |
| 50 |
| 60 |
| 70 |
| 80 |
| 90 |
| 100 |
| Original model |
| p-value |
| 0.77869 |
| 0.90639 |
| 0.90014 |
| 0.42233 |
| 0.61180 |
| 0.99849 |
| Doc2Vec (Sh) |
| p-value |
| 0.95348 |
| 0.98867 |
| 0.99588 |
| 0.97935 |
| 0.99791 |
| 0.88989 |
| Doc2Vec (Sh+R) |
| p-value |
| 0.91981 |
| 0.99952 |
| 0.80851 |
| 0.99314 |
| 0.97609 |
| 0.93013 |
| MiniLM (Sh) |
| p-value |
| 0.91337 |
| 0.99382 |
| 0.65138 |
| 0.84190 |
| 0.45006 |
| 0.94460 |
| NPNet (Sh) |
| p-value |
| 0.80502 |
| 0.90611 |
| 0.99869 |
| 0.75723 |
| 0.49358 |
| 0.96711 |
In the first case, we compare the original model, which shows a decreasing linear trend as the training data size increases (black dashed line). However, its confidence interval also narrows after considering 80% of the training data, suggesting reduced variability around the MSE mean and greater confidence that the true MSE value lies within this (black) interval. Comparing all intervals, this would indicate possible limitations in further reducing the error for the original model. When examining the curves for the cases where Doc2Vec is used to encode either only the insert shapes (in blue) or both the insert shapes and the recipe (in green), we observe that the average MSE follows a similar decreasing trend. Additionally, we see a narrowing of their respective confidence intervals after considering 70% of the data, and these intervals cover a slightly lower range than the original model’s, suggesting that further reduction in error may be possible, as the bounds of both confidence intervals (blue and green) are slightly below that of the original model. This could also be achieved by incorporating more data or by fine-tuning the model’s parameters.
In the other two cases, corresponding to the application of all-mpnet-base-v2 and all-MiniLM-L12-v2 transformers, we observe similar behavior to Doc2Vec, except that the average error in each case is very close to the average MSE of the original model, indicating no significant improvements. Furthermore, in the case of transformers, considering 80% of the training data results in a widening of their confidence intervals, suggesting that these representations require as much data as possible to achieve performance equal to or only marginally better than that of the original model. This approach could reduce the variability of the MSE mean, resulting in narrower intervals and providing greater confidence that the true value falls within this range.
5.5 score
Similarly to the previous section, we will analyze the evolution of the metric as a function of the proportion of data used for training, through -fold cross-validation.
To assess whether the obtained values follow a normal distribution, we again use the Kolmogorov-Smirnov test. As shown in Table 5, the p-values (> 0.05) support the assumption of normality.
| Fraction |
| of data (%) |
| 50 |
| 60 |
| 70 |
| 80 |
| 90 |
| 100 |
| Original model |
| p-value |
| 0.53053 |
| 0.30654 |
| 0.49133 |
| 0.57835 |
| 0.96425 |
| 0.74970 |
| Doc2Vec (Sh) |
| p-value |
| 0.51809 |
| 0.83924 |
| 0.41014 |
| 0.47981 |
| 0.99058 |
| 0.83713 |
| Doc2Vec (Sh+R) |
| p-value |
| 0.74102 |
| 0.92346 |
| 0.41024 |
| 0.62140 |
| 0.98585 |
| 0.71981 |
| MiniLM (Sh) |
| p-value |
| 0.72533 |
| 0.83543 |
| 0.32626 |
| 0.42428 |
| 0.97045 |
| 0.74681 |
| NPNet (Sh) |
| p-value |
| 0.76436 |
| 0.69900 |
| 0.47973 |
| 0.49954 |
| 0.98269 |
| 0.76401 |
Then, in the first case (see Fig. 8), we explore the original model, which shows an increasing trend as the training data size grows (dashed black line). However, its confidence interval also widens downward once 90% of the training data is considered, which means it limits the potential for improvement in this metric, as the bounds of the other (blue and green) confidence intervals within which the true value of is likely to fall are slightly above those of the original model.
In contrast, when examining the curves for models using Doc2Vec to encode only the shapes of the inserts (in blue) or both the shapes of the inserts and the recipe (in green), the average follows a similar increasing trend, slightly above the black curve. Additionally, their respective confidence intervals also widen once 90% of the data is included.
In the remaining two cases (see Fig. 8), similar behaviors are observed, although the average scores in each case are closer to the original model’s average , indicating no significant improvements. For transformers, the same widening of confidence intervals is noted when considering 90% of the training data, suggesting that the model requires as much data as possible to achieve only slight improvements in performance. Again, this approach helps to reduce the variability of the MSE mean, ensuring that the true value falls within this interval.
Furthermore, parameter optimization was performed on the XGBOOST regression parameters for each of the models used for encoding categorical variables. In all cases, the optimized parameters were found to be close to those originally set. Then, a similar analysis, including 10-fold cross-validation with varying portions of the data for training, yielded results consistent with those presented in the previous sections.
Finally, to analyze the performance of the XGBOOST model when varying the dimensionality of the latent spaces generated by transformers and embedding, we also estimated the confidence intervals for the average MSE and . Considering much larger latent space dimensions for the Doc2Vec model do not show significant improvements compared to those already achieved with a dimension of 3. However, it is more interesting to observe the effects when reducing the latent space dimension of the transformers to 3 (cf. Appendix A, Fig. 15 and Fig. 16). The main advantages here are a slight improvement in metrics and increased efficiency of the predictive model, as it has to handle fewer variables.
5.6 Feature importance and Shapley Analysis
Recall that for the feature importance with total_gain is based on the total reduction in error achieved by splitting all the trees, indicating how much each input feature contributes to reducing the overall error of the XGBOOST model. Similarly, the Shapley analysis focuses on each feature contribution to the model predictions. In this context, we will analyze how the implementation of transformers for the encoding of categorical features affects the corresponding importance scores for the input features.
For the original model (see Fig. 9), which uses binary encoding for handling categorical features, we observe that numerical features (in blue) are the most influential in terms of feature importance scores, particularly those related to total area and surface variability. In contrast, categorical variables related to insert geometries (in yellow) and the recipe (in green) have considerably lower scores. Shapley analysis shows a similar trend where numerical features have higher scores than categorical ones. This suggests that potentially critical inputs for the process are neglected as categorical variables.
Both feature importance and Shapley analysis suggest that, although some numerical features remain the most important, we observe a slight increase in the importance of insert shapes after being encoded with Doc2Vec (see Fig. 10). It is also important to note that the number of input features has tripled following preprocessing and encoding. This indicates that incorporating more useful information into the regression model is indeed possible by transforming the descriptions of insert shapes using an embedding.
Finally, when we use more sophisticated models as transformers to encode categorical features (see Figs. 11 and 12), a similar trend continues in terms of feature importance and Shapley analysis: categorical variables related to the shapes of the inserts are ranked within the top 5. This further emphasizes the value of having interpreted textual descriptions more robustly for inclusion in the regression model.
6 Conclusions
This case study highlights the significant benefits derived from leveraging meaningful embeddings, particularly through the use of transformer models, for processing categorical variables. Transformers, known for their ability to capture intricate patterns and relationships in data, enable categorical variables to be represented in a manner that captures their semantic and contextual meaning. By doing so, the embeddings generated reflect not only the distinct values of these categorical variables but also the underlying relationships between them, which might otherwise be missed by traditional encoding techniques such as one-hot encoding.
The ability of transformers to handle categorical variables more effectively results in a substantial improvement in the performance and efficiency of predictive models. By embedding categorical data in a way that is contextually rich and semantically meaningful, the model can extract deeper insights and make more accurate predictions. This approach ensures that models are better equipped to handle complex, real-world datasets where categorical variables often play a crucial role in determining outcomes.
In addition to enhancing predictive capabilities, this method improves the assessment of critical process parameters. Categorical variables can now be considered in a manner that is more representative of actual observations and more aligned with the way data behaves in real-world scenarios. This leads to a more nuanced understanding of the interactions between variables and offers a more comprehensive approach to process optimization. By embedding categorical data in a meaningful way, models can incorporate these variables into their structure more naturally, enhancing both the interpretability and the accuracy of the results.
Looking forward, there is great potential in expanding this work into sensitivity analysis and uncertainty quantification. These are critical areas for model evaluation, as they provide insights into how changes in input variables affect the output and help quantify the uncertainty associated with predictions. Future research could focus on exploring more advanced machine learning methods and alternative mathematical frameworks to further improve both the accuracy and interpretability of models. Methods such as Bayesian neural networks, variational inference, and probabilistic graphical models could provide robust frameworks for uncertainty quantification, offering deeper insights into the reliability of predictions.
Moreover, by integrating these advanced techniques into the existing framework for processing categorical variables, future studies could refine sensitivity analysis, making it more adaptable to complex, high-dimensional data. This would not only boost the performance of models but also provide more granular control over how categorical data influences predictions, enabling more informed decision-making in critical applications.
Acknowledgments
This research was funded in part by the Luxembourg National Research Fund (FNR), grant reference [16758846]. For the purpose of open access, the authors have applied a Creative Commons Attribution 4.0 International (CC BY 4.0) license to any Author Accepted Manuscript version arising from this submission. PP gratefully acknowledges funding from the FSTM in the University of Luxembourg.
Appendix A Appendix
A.1 Similarity after dimensionality Reduction
A.2 XGBOOST performance after dimensionality Reduction